OpenStudy (anonymous):

i need help!!!! 2x over x-1 minus 7 over x-2

OpenStudy (anonymous):

2x/ (x-1) - 7/(x-2) = (2x(x-2) - 7 (x-1) )/ (x-2)(x-1) = (2x^2-4x - 7x +7) / (x-2)(x-1) = 2x^2 -11x +7 / (x-2)(x-1)

OpenStudy (anonymous):

Thank You!